Find the average horsepower, P =? required to raise a 150 kg drum to a height of 20 m in a time of 1 minute. SOLUTION: P = wk/t = mgH/t

1
Expert's answer
2022-02-21T12:05:55-0500

Given:

"m=150\:\rm kg"

"h=20\:\rm m"

"t=1\rm \: min=60\: s"


The average power

"P=\frac{W}{t}=\frac{mgh}{t}"

"P=\rm\frac{150*9.8*20}{60}=490\:W=0.657\; hp"